OpenStudy (anonymous):

tan^2x = cos(150)tan(-330)/sin(150)

OpenStudy (ash2326):

We have \[ \tan ^2x= \cos(150) \times \frac{\tan(-330)}{\sin 150}\] @meow.fz do you know the value of cos 150, sin 150 and tan (-330)\

OpenStudy (anonymous):

cos(150) = -\[\sqrt{3}\div2\] sin(150) = \[\sqrt{3}\div2\] tan(-330) = \[1\div \sqrt{3}\]

OpenStudy (ash2326):

great but sin 150 is [\frac{1}{2}\] no worries, Let's proceed with the problem, let's substitute the values \[\tan^2 x= \frac{-\sqrt{3}}{2}\frac{ \frac{1}{\sqrt 3}}{\frac{1}{2}} \] now let's cancel the common terms \[\tan^2 x= \frac{-\cancel {\sqrt{3}}}{\cancel2}\frac{ \frac{1}{\cancel{\sqrt 3}}}{\frac{1} {\cancel2}} \] so we get \[ \tan^2 x=-1\] \[=> \tan x= \pm i( i=\sqrt{-1})\] so no real solution

OpenStudy (ash2326):

\[\sin 150=\frac{1}{2}\]

OpenStudy (ash2326):

Did you understand?

OpenStudy (anonymous):

okay cool ,thank you sooo much dude , i really appreciate it , i have this assignment due tomorrow , and im battling with putting the answers together , thanks , i understood :)

OpenStudy (ash2326):

welcome meow:D and Welcome to Open Study:D

OpenStudy (anonymous):

thank you :)